OpenStudy (anonymous):

What is the slope of the line passing through the points (4, -2) and (-5, 4)?

OpenStudy (ipwnbunnies):

Use the formula for slope for two points: \[slope = \frac{Y2 - Y1}{X2 - X1}\] Where (X1,Y1) is one point, and (X2,Y2) is the other point.

OpenStudy (anonymous):

2/3 3/2 -2/3 -3/2

OpenStudy (anonymous):

I think it's d

OpenStudy (ipwnbunnies):

Did you do the math. >.>

OpenStudy (ipwnbunnies):

\[(X1,Y1) = (4,-2)\] \[(X2,Y2) = (-5, 4)\]

OpenStudy (ipwnbunnies):

\[slope = \frac{4 - (-2)}{-5 - 4}\]

OpenStudy (anonymous):

I got 6/-9

OpenStudy (anonymous):

My answer is not in the answer choice

OpenStudy (ipwnbunnies):

You have to simplify the fraction. 6 and 9 are both divisible by 3.

OpenStudy (anonymous):

-2/3

OpenStudy (ipwnbunnies):

Yesh. :3